OpenStudy (logan13):

What is 3 root of 250? A. 5 3 root of 2 B. 2 3 root of 5 C. 5 3 root of 3 D. 3 3 root of 5

OpenStudy (logan13):

@Mehek14

OpenStudy (logan13):

@TheSmartOne

OpenStudy (logan13):

C

TheSmartOne (thesmartone):

no

OpenStudy (logan13):

Hmm

TheSmartOne (thesmartone):

Hint: \(\Large\bf \sqrt[3]{250} = \sqrt[3]{5^3 \times 2}\)

OpenStudy (phi):

you should notice that 250 is 5*5*5 * 2 if you have a *triple* inside a cube root (i.e. 3rd root), you can "pull out the triple" and put just one of the 5's out front. the 2 has to stay inside

OpenStudy (ssddss02):

Your hint is the answer lol @TheSmartOne look at the end of the hint eqtion he gave you it will give you the answer

OpenStudy (ssddss02):

If you understand how where doing it I can give you the answer @logan13

OpenStudy (ssddss02):

Its A. your welcome.

OpenStudy (ssddss02):

I would love a medal